OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. The sum of a positive number and its square is 20. What is the number? A. 1 B. 2 C. 3 D. 4 E. 5

OpenStudy (bahrom7893):

x+x^2 = 20 x+x^2-20 = 0 solve for x

OpenStudy (anonymous):

x+x^2=20 x^2+x-20=0 (x+5)(x-4)=0 You want the positive value, ignore the negative

OpenStudy (bahrom7893):

x = 4

OpenStudy (anonymous):

I got tht answer too

OpenStudy (anonymous):

thnk u guys